# Deep learning modelling of manufacturing and build variations on multi-stage axial compressors aerodynamics

### v3

@@ 4 Conclusion
This study presents the development and application of a novel deep learning framework for real-time predictions of the impact of tip clearance variations on the flow field and overall performance of multi-stage axial compressors. The proposed C(NN)FD architecture achieves real-time accuracy comparable to the CFD benchmark. Our methodology is generalisable because of the flow field predictions, which are utilized to calculate using the relevant physical equations the corresponding radial profiles, 1D averages, stage-wise distributions and overall performance. Our methodology is demonstrated to be scalable to industrial applications because of physics-based selective used of relevant parts of the CFD solution. The deployed model, is readily integrated within the manufacturing and build process of gas turbines, thus providing the opportunity to analytically assess the impact on performance and to take the appropriate measures to potentially avoid repeating expensive physical tests and $CO_{2}$ emissions.

### v4

@@ 4 Conclusion
This paper demonstrates the development and application of a deep learning framework for predictions of the flow field and aerodynamic performance of multi-stage axial compressors. A physics-based dimensionality reduction unlocks the potential to use supervised learning approaches for flow-field predictions, as it re-formulates the regression problem from an un-structured to a structured one, as well as reducing the number of degrees of freedom. The proposed architecture also has a competitive advantage compared to traditional ”black-box” surrogate models, as it provides explainability to the predictions of variations in overall performance by identifying the corresponding aerodynamic drivers. This is applied to model the effect of manufacturing and build variations, as the associated scatter in efficiency is known to have a significant impact on the overall performance and $CO_{2}$ emissions of the gas turbine, therefore posing a challenge of great industrial and environmental relevance. The proposed architecture is proven to achieve an accuracy comparable to that of the CFD benchmark, in real-time, for an industrially relevant application. The deployed model, is readily integrated within the manufacturing and build process of gas turbines, thus providing the opportunity to analytically assess the impact on performance with actionable and explainable data.